When incorporating, it is important to consider whether or not the company’s business deals internationally. 1. Summarize the impact of foreign exchange rates on the company’s financial statements. What risks do foreign exchange rates pose? 2. What are the two methods used to translate financial statements and how does the functional currency play a role in determining which method is used? 3. Compose a hypothetical example to demonstrate the translation process using the two methods. Ensure all information is entered accurately.
Answer #1

1)Companies with international deals should translate the transactions from their functional currency to a single reporting currency as required by the accounting guidelines. Significant shifts in currency exchangerates can have a dramatic effect on the company’s revenue and profit trends that means fluctuation in exchange rates can affect financial statement trends.The foreign exchange gain or loss will reflect as an income or expenses in the income statement and the same will be correspondingly reflected in balance sheet as well.However the gain or loss from the exchange rate can have either a positive or negative effect depending on which currency the transaction is designated in and which direction the exchange rate moves. Foreign currency risk or Transaction risk is the risk associated with Foreign Exchange Rate.

2) Current rate method (or closingrate) and the temporal method are the two methods used to translate financial statements .Primary currency of a foreign entity’s operating environment is its functional currency. The functional currency orientation results in the following rule:

Functional currency Translation method Translation adjustment US DOLLAR Temporal method Gain(loss) in Net income Seperate com

3)Suppose P&G has a subsidiary in Thailad. Its subsidiary in Thailand reported the following:

(All in the local currency i.e. Thai baht): Current assets = 5,000,000. Fixed assets = 35,000,000. Total assets = 40,000,000. Current liabilities = 3,000,000. Long term debt = 12,000,000. Equity = 25,000,000. Total liability and equity = 40,000,000.

Now the current exchange rate will translate all the assets and liabilities (which are in Thai baht) to USD. The current exchange rate will be used. Current rate = 1 baht = 0.028 USD. The equity is translated using the rate applicable on the date of issuance. Suppose the rate on issuance date was 1 baht = 0.025 USD. Any balance will be shown as cumulative translation djustment (CTA).

So, current assets in USD will be = 5,000,000*0.028 = $140,000. Fixed assets = 35,000,000*0.028 = $980,000. Total assets = $1,120,000. Current liability =  3,000,000*0.028 = $84,000. Long term debt = 12,000,000*0.028 = $336,000. Equity = 25,000,000*0.025 = $625,000. Total liability and equity = $1,045,000

Difference = 1,120,000 - 1,045,000 = $75,000 (assets - liabilities). This will be shown as CTA in the balance sheet as the adjusting amount.

Assets Liabilities
Current assets 140,000.00 Current liabilities 84,000.00
Fixed assets 980,000.00 Long term debt 336,000.00
Equity 625,000.00
CTA 75,000.00
Total assets 1,120,000.00 Total liability and equity 1,120,000.00

Temporal rate translation method  uses exchange rates based on the time assets and liabilities are acquired or incurred.

Current assets are translated using current exchange rate, fixed assets are translated using historical exchange rate, equity is translated using the rate at time of issuance.

Assets in baht Relevant rate Description Liabilities in baht Relevant rate Description
Current assets 5,000,000.00 0.028 Current rate Current liabilities 3,000,000.00 0.028 Current rate
Fixed assets 35,000,000.00 0.026 Historical rate Long term debt 12,000,000.00 0.027 Historical rate
Equity 25,000,000.00 0.025 Rate on date of issuance
Total assets 40,000,000.00 Total liability and equity 40,000,000.00

Using the given rates, conversions are done to obtain the numbers in USD:

Assets in USD Liabilities in USD
Current assets 140,000.00 Current liabilities 84,000.00
Fixed assets 910,000.00 Long term debt 324,000.00
Equity 625,000.00
CTA 17,000.00
Total assets 1,050,000.00 1,050,000.00